OpenStudy (anonymous):

For 1b, I have the alphas for initial, midpoint, and final configuration. How do I calculate the spline from this point? The notes and hw solutions do not make sense to me.

OpenStudy (anonymous):

From the notes, \[P_{simp}*\left(\begin{matrix}\alpha_i \\ \alpha_v\\ \alpha_f\end{matrix}\right)=\left(\begin{matrix}a_0 \\ a_1\\a_2\\a_3\\b_0\\b_1\\b_2\\b_3\end{matrix}\right)\]

OpenStudy (anonymous):

So Psimp = (alphai; alphav; alphaf)^-1 * (a0;a1;a2;a3;b0;b1;b2;b3) Is that what we are solving for?

OpenStudy (anonymous):

No we want the coefficients for our two splines: a0-b3. There is a formula for Psimp in the notes. It only depends on the quantity t which in this case is 2.5.

OpenStudy (anonymous):

a0 = pi ; a1 = 0 ; a2 = 3/(tf)^2 * (pf -pi) ; a3 = 2/(tf)^3 * (pi-pf) Is this correct? If so, what is pi and pf?

OpenStudy (anonymous):

This would be correct if we were only trying to connect two points with a spline. Here we are trying to connect three points with 2 splines. For each joint pi = alpha(0), pf = alpha(5), and pv = alpha(2.5). You need one trajectory that goes from 0 to 2.5(accounted for by the as) and one that goes from 2.5 to 5(accounted for by the bs).

OpenStudy (anonymous):

Ok, so using the equations I just posted, I can calculate a0-a3 for my two splines. (giving me p(t) for both splines). For q(t): I don't see anything conclusive on how to calculate b0-b3

OpenStudy (anonymous):

p(t) is the function for 0<t<tv and q(t-tv) is the function for 2.5<t<tf. The cannot be calculcated separately because we want their velocities and accelerations to match at t=2.5 All the a's and b's are calculated at once using the formula I gave you above.

OpenStudy (anonymous):

All I see that is relevant in the notes is : "8 coefficients, 8 equations": p(0) = alphai q(0) = alphav p(t0) = a0 qdot(tf-tv) = 0 pdot(0) = 0 q(tf-tv) = alphaf pdot(tv) = qdot(v) pdoubledot(tv) = qdoubledot(0)

OpenStudy (anonymous):

I made two columns of four equations here

OpenStudy (anonymous):

Yes that's part of the way we derived Psimp. I think if you look in your notes you will find an 8x3 matrix that depends on t. If you multiply that matrix by [alphai; alphav; alphaf], you will get all the 8 coefficients you need.

OpenStudy (anonymous):

ok, i found that. What do we use as our t value for our two different splines?

OpenStudy (anonymous):

There is only 1 t value, it's the time step and is 2.5 in this case.

OpenStudy (anonymous):

So if I plug in 2.5 for in the 3x8 matrix and multiply it by my alpha matrix, I will get my final answer?

OpenStudy (anonymous):

for t*

OpenStudy (anonymous):

What do you mean by alpha matrix?

OpenStudy (anonymous):

(alphai, alphav, alphaf)^T

OpenStudy (anonymous):

Yes for each joint. So in total you will have three sets of a's and b's

OpenStudy (anonymous):

So for my final answer, I will simply have three sets of a's and b's?

OpenStudy (anonymous):

Yes. for each joint you need two splines: one that goes to 2.5s and the other goes 2.5-5s

OpenStudy (anonymous):

So, a p(t) function and q(t) functin for each joint?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

how do we get our alphav values from our calculated g(2.5)?

OpenStudy (anonymous):

inverse kinematics

OpenStudy (anonymous):

the triangle method?

OpenStudy (anonymous):

Yeah it uses triangle geometry. You did it 2 homework ago. You can check the solutions if you don't remember.